# Rational Design Heterobilayers Photocatalysts for Efficient Water Splitting Based on 2D Transition-Metal Dichalcogenide and Their Janus

### v1

@@ Abstract
Direct Z-scheme heterostructures with enhanced redox potential are increasingly regarded as promising materials for solar-driven water splitting. This potential arises from the synergistic interaction between the intrinsic dipoles in Janus materials and the interfacial electric fields across the layers. In this study, we explore the photocatalytic potential of 20 two-dimensional (2D) Janus transition metal dichalcogenide (TMDC) heterobilayers for efficient water splitting. Utilizing density functional theory (DFT) calculations, we first screen these materials based on key properties such as band gaps and the magnitude of intrinsic electric fields to identify promising candidates. We then evaluate additional critical factors, including carrier mobility and surface chemical reactions, to fully assess their performance. The intrinsic dipole moments in Janus materials generate built-in electric fields that enhance charge separation and reduce carrier recombination, thereby improving photocatalytic efficiency. Furthermore, we employ the Fröhlich interaction model to quantify the mobility contributions from the longitudinal optical phonon mode, providing detailed insights into how carrier mobility, influenced by phonon scattering, affects photocatalytic performance. Our results reveal that several Janus-TMDC heterobilayers, including WSe2-SWSe, WSe2-TeWSe, and WS2-SMoSe, exhibit strong absorption in the visible spectrum and achieve solar-to-hydrogen (STH) conversion efficiencies of up to 33.24%. These findings demonstrate the potential of Janus-based Z-scheme systems to overcome existing limitations in photocatalytic water splitting by optimizing the electronic and structural properties of 2D materials. This research highlights a viable pathway for advancing clean energy generation through enhanced photocatalytic processes.
@@ 1 Introduction
Due to escalating environmental pollution and energy challenges, the search for alternative energy sources has gained significant attention in the modern era. Hydrogen energy is increasingly regarded as an innovative technology for the future, providing the dual advantages of being both clean and renewable. 1, 2, 3. Photocatalytic water splitting has garnered significant attention in global scientific and industrial communities, which convert solar energy through the splitting process of water to H2 energy. This utilization of solar light energy is accelerating the carbon neutralization process in the world 4. The photocatalytic technique is based on generating charge carriers within a photocatalyst system under suitable light irradiation. Therefore, the primary challenge is discovering a highly efficient photocatalyst that exhibits favorable photoelectronic and optical properties for STH energy conversion. 5, 6. Maximizing the efficiency of photocatalytic hydrogen generation depends on the control optimization of the critical parameters, including i) an optimal band gap and appropriately aligned conduction and valence band of the photocatalyst to effectively harvest visible light and facilitate the redox reactions required for water splitting; ii) a high surface area and well-engineered morphology, alongside a crystalline structure, enhancing light absorption and efficiently promoting charge mobility; iii) suppression of electron-hole recombination and enhanced charge carrier mobility for efficient charge transport to reactive sites; iv) reaction conditions, including pH, temperature, and the incorporation of sacrificial agents 6, 5, 7, 8.
@@
Among the various photocatalysts, two-dimensional (2D) materials offer many favorable properties, such as large surface area, quantum confinement effects, strong light absorption enhanced charge separation, and structural flexibility 9, 10, 11, 12. One of the standout features of 2D materials is their planar structure, which offers a high density of surface atoms and ample space for integration with co-catalysts. This results in more available active sites for catalytic reactions, and the potential creates new active sites, enhancing overall catalytic efficiency 13. Additionally, the significantly reduced thickness of 2D materials enables charge carriers to travel shorter distances to reach the surface. This leads to decreasing electron-hole recombination and preserves more charge carriers for surface reactions. These unique characteristics make 2D materials promising candidates for improving the performance of photocatalytic systems 13, in which TMDC materials have been considered as the most potential applicant due to possessing efficient visible light absorption and high charge carrier mobility enhancing charge separation. The TMDCs have demonstrated good chemical stability and robust light-matter interactions, further improving their efficacy in light-driven processes. Despite these strengths, TMDCs are facing plausible limitations in photocatalytic hydrogen generation due to high exciton recombination and strong electrostatic coulomb interactions, leading to inefficient photogenerated charge carriers 14, 15. In comparison with conventional TMDCs, Janus TMDCs include different chalcogen atoms, breaking the out-of-plane symmetry 16. Thus, the Janus TMDCs will create out-of-plane dipoles between the top and bottom layers that could generate an intrinsic electric field and enhance the photocatalytic performance in Janus-TMDCs 17, 18, 19.
@@
Based on the above reasons, the heterobilayers combining TMDCs and their Janus materials that have been successfully synthesized 20, 21 could benefit photocatalytic water splitting. In monolayer 2D material, both the hydrogen evolution reaction (HER) and the oxidation evolution reaction (OER) occur on the same surface of the 2D material, resulting in a higher recombination rate of photogenerated charge carriers 6, 5. A promising solution to this problem is the van der Waals (vdW) heterobilayers of two semiconductor layers. In particular, the valence band maximum (VBM) and conduction band minimum (CBM) are integrated into both semiconductor sides, which are known as type II and Z-scheme heterostructure systems. A distinct behavior of the Z-scheme and type-II heterostructures is their unique electronic structure. Notably, the Z-scheme junction features a suitable band alignment and a wide directional built-in electric field. As a result, photogenerated charge carriers with a large redox potential are well-preserved and spatially separated within the Z-scheme heterostructures. Combining two narrow band gap semiconductors into a Z-scheme system could solve the challenge of a single platform or type-II junctions, where strong redox ability and light absorption region may not be obtained simultaneously 22, 23. This configuration allows HER and OER activity on separate surfaces. On the other hand, the heterobilayer with the Janus structure generates an intrinsic electric field, which not only enhances the photocatalytic performance but also accelerates the reaction kinetics by forming synergistic surface properties that allow selective adsorption of water molecules on each surface 24, 25, 26, 27. Moreover, the configuration of the heterobilayer can determine whether photogenerated carriers follow a Z-scheme or a traditional type-II pathway 27. However, the diversity of possible Janus-TMDC combinations in heterobilayers poses a significant challenge in determining the optimal configuration for photocatalytic water splitting. In this study, we provide a theoretical framework for rapidly identifying effective Janus-TMDC pairs through a top-down approach: candidates are initially filtered based on band alignment, and further refinement follows as needed. Another key challenge is the intrinsic polarization of Janus materials, which uniquely affects their electrical mobility. However, this factor is often neglected in previous Janus or TMDC studies for photocatalytic 28, 1, 29. To address this, we have adopted Fröhlich interaction model 30, 28 to account for the mobility adjustments in Janus heterostructures accurately.
@@ 2 Theoretical Approach and Computational Details
In this research, we employed DFT for calculations using the Quantum ESPRESSO package 31, 32, 33. The generalized gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional was utilized as implemented in the standard solid-state pseudopotentials (SSSP) library 34, 35, 36. Van der Waals interactions were incorporated using the Grimme DFT-D2 method 37, 38, 39. To address the underestimation by PBE in predicting band edge positions and band gaps, we used the Heyd-Scuseria-Ernzerhof (HSE) hybrid-functional with a mixing parameter 25% 40, 41. The kinetic cutoff energy for wavefunctions ($E_{\text{cut}}$) was set to 800 eV, and the Monkhorst-Pack k-point grid was set to $12\times 12\times 1$ for PBE and $6\times 6\times 1$ for HSE based on the convergence test. All structures were fully relaxed using the BFGS scheme 42, with convergence criteria for force and energy differences set to $1\times 10^{-4}$ eV/Å and $1\times 10^{-8}$ eV, respectively. To prevent interlayer interactions, a relatively large vacuum layer of 30 Å was considered along the $z$-direction. To examine the Gibbs free energy and the mobility of the hydrogen atom, we adopted a $3\times 3\times 1$ supercell with 54 atoms. For the carrier mobility calculation, a rectangle $\sqrt{3}\times 1\times 1$ supercell was used to obtain the effective mass of carriers along $x$- and $y$-directions. The Wannier90 code and the VESTA software were used for pre-processing and post-processing 43, 44.
@@ 3.1 Geometric and Electronic Structures
The Janus-TMDC heterobilayers have many configurations by selecting the atoms at the interface between two layers. Replacement of these atoms results in changes in the difference in electronegativity between layers, which, in turn, modifies their optoelectrical properties and influences the photocatalytic capabilities of the material. Based on this concept, Janus MoSSe, WSSe, and their heterobilayers with TMDC, such as MoSSe-MoS2, have already been successfully synthesized 20, 21. Following these achievements, recent theoretical research has focused on other Janus-type materials based on molybdenum (Mo) and tungsten (W) 45, 46, 47, 48, 49, 50. Thus, we selected Mo and W as the metals of interest because of their promising possibilities for future synthesis and applications in water splitting. For the chalcogenides, we used combinations of sulfur (S), selenium (Se), and tellurium (Te), resulting in the formation of 2D vdW structures comprising conventional TMDC and Janus materials.
Figure 1: Visual representations of the vdW Janus-TMDC structure and its Brillouin zone: (a) top view; (b) side view; (c) Brillouin zone indicating high-symmetry points.
The geometric structures of these 20 materials were designed based on the previously synthesized AA stacking mode of the MoSSe-MoS2 structure, systematically replacing the metal and chalcogenide atoms, as shown in Figure 1. The optimized lattice parameters are listed in Table 1. We note that the structures of MoS2-TeMoSe and WS2-TeWSe did not converge during the optimization calculations, indicating that these configurations are unstable. This instability can be attributed to the significant differences in electronegativity between the S and Te atoms.
@@ 3.2 Photocatalytic characteristic
After completion of structural optimizations, we evaluated the potential of 18 materials as photocatalysts (see Table S1 in Supporting Information). To determine the electronic properties critical for photocatalytic activity, we performed partial energy band structure for 18 material combinations, in which the partial band structures show the contributions of individual layers in the heterostructures to the electronic states near the band edges (see Figure S1, Supporting Information). By projecting the electronic states onto the atomic orbitals of the Janus and TMDC layers, we identified which layers predominantly contribute to the VBM and CBM. This approach allowed us to classify the materials into Type I band alignment, where both the VBM and CBM are contributed by the same layer, or Type II alignment and Z-scheme alignment, where they are contributed by different layers, based on the spatial localization of their VBM and CBM.
@@
To be considered as a photocatalyst for water splitting, the theoretical band gap should be larger than 1.23 eV to overcome the reduction and oxidation potential of H2O 1, 51. However, due to the difference in electrostatic potential between the layers contributing to the VBM and CBM, even materials with smaller band gaps can act as photocatalysts. As shown in Table S1 (Supporting Information), the materials with Type I band alignment, including MoSe2-SeMoTe, WSe2-SeWS, and WSe2-SeWTe, have band gaps less than 1.23 eV. Thus, they are unsuitable for water-splitting applications.
@@
For the remaining materials, we calculated the electrostatic potential curves, as shown in Figure S2 in Supporting Information. The calculations include a dipole correction to suppress artificial dipole fields by introducing an additional ramp-shaped potential 42. The intrinsic electric field generated by the Janus layer in the heterostructure results in a significant potential difference between the two layers, leading to a difference in work function ($\Delta\Phi$) as calculated and presented in Table 1. Note that $\Delta\Phi$ is determined by taking the difference between the work functions of the TMDC and Janus layers, following this equation 14:
$\Delta\Phi=V_{\text{TMDC}}-V_{\text{Janus}}.$ (1)
@@
The band alignment plays a critical role in determining the photocatalytic properties of Janus-TMDC heterobilayers. When the work function difference between the layers, denoted as $\Delta\Phi$, is positive ($\Delta\Phi>0$), this indicates that the vacuum level of the TMDC layer is higher than that of the Janus layer. In such a configuration, the TMDC layer predominantly contributes to the VBM, while the Janus layer largely influences the CBM. This arrangement leads to a Z-scheme band alignment, which effectively broadens the separation between the CBM and VBM, enhancing charge carrier separation and improving photocatalytic activity (e.g., WSe2-SWSe and WS2-SMoSe, as shown in Figure 2). Conversely, when $\Delta\Phi$ is negative ($\Delta\Phi<0$), meaning the vacuum level of the Janus layer is higher than that of the TMDC layer, the Janus layer mainly contributes to the VBM, and the TMDC layer to the CBM. This configuration also results in a Z-scheme alignment (e.g., WSe2-TeWSe, as shown in Figure 2).
@@
In cases where $\Delta\Phi$ is negligible or does not meet the conditions above, the system exhibits a type II band alignment. The effective band gap in a Z-scheme configuration can be expressed as $E_{\text{eff}}=E_{g}+\Delta\Phi$, where $E_{g}$ represents the intrinsic band gap, and $\Delta\Phi$ reflects the built-in potential difference at the heterostructure interface. Materials with $E_{\text{eff}}>1.23$ eV are considered promising candidates for photocatalytic applications, and their performance is further evaluated through calculations of STH efficiency. The $E_{\text{eff}}$ values for the materials studied are summarized in Table 1.
Table 1: Calculated band gap energies, work function differences, and effective band gaps for various Janus-TMDC heterobilayers.
[…]
MoS2-SeMoTe 0.315 0.82 II -
MoS2-TeMoSe - - - -
MoSe2-SeMoTe 1.044 0.84 I 1.044
MoSe2-TeMoSe 0.481 -0.6 Z-scheme 1.081
MoS2-SMoSe 1.313 0.82 I 1.313
MoS2-SeMoS 0.68 -0.64 Z-scheme 1.32
MoSe2-SMoSe 0.494 0.71 Z-scheme 1.204
MoSe2-SeMoS $1$0.73 I 1.111
WS2-SWSe 1.437 0.79 I 1.437
WS2-SeWS 0.815 -0.66 Z-scheme 1.475
WSe2-SWSe 0.604 0.69 Z-scheme 1.294
WSe2-SeWS 1.211 -0.7 I 1.211
WS2-SeWTe 0.411 0.93 II -
WS2-TeWSe - - - -
WSe2-SeWTe 1.103 0.8 I 1.103
WSe2-TeWSe 0.969 -0.63 Z-scheme 1.599
MoS2-SWSe 1.562 0.83 II 0.732
MoS2-SeWS 0.952 -0.58 Z-scheme 1.532
WS2-SMoSe 0.937 0.79 Z-scheme 1.717
WS2-SeMoS 1.517 -0.71 Z-scheme 2.227

### v2

@@ Abstract
Direct Z-scheme heterobilayers with enhanced redox potential are viewed as promising for solar-driven water splitting, arising from the synergy between intrinsic dipoles in Janus materials and interfacial electric fields across the layers. This study explores 20 two-dimensional Janus transition-metal dichalcogenide (TMDC) heterobilayers for efficient water splitting. Using density-functional theory (DFT) calculations, we screen them based on band gaps and intrinsic electric fields to identify promising candidates, then further assess carrier mobility and surface chemistry to fully evaluate their overall performance. By examining the alignment of synthetic and internal electric fields, we distinguish between Type-I, Type-II, and Z-scheme configurations, enabling the targeted design of optimal photocatalytic materials. Furthermore, we employ the Fröhlich interaction model to quantify the mobility contributions from the longitudinal optical phonon mode, providing detailed insights into how carrier mobility, influenced by phonon scattering, affects photocatalytic performance. Our findings demonstrate the potential of Janus-based Z-scheme systems to overcome existing limitations in photocatalytic water splitting by optimizing the electronic and structural properties of 2D materials, highlighting a viable pathway for advancing clean energy generation through enhanced photocatalytic processes.
@@ 1 Introduction
Photocatalytic water splitting that converts solar energy through the splitting process of water to hydrogen H2 energy has garnered significant attention in global scientific and industrial communities 1, 2, 3. Among the various photocatalysts, 2D materials stand out for water splitting due to their large surface area, quantum confinement, strong light absorption, enhanced charge separation, and structural flexibility 4, 5. Their planar structure features a high density of surface atoms and ample space for co-catalyst integration, boosting available active sites and overall efficiency 6. Moreover, the reduced thickness of 2D materials shortens charge-carrier travel distances to the surface, mitigating electron-hole recombination and preserving more carriers for catalytic reactions. These characteristics make 2D materials promising photocatalysts, particularly TMDCs, which benefit from efficient visible light absorption and high charge-carrier mobility for enhanced charge separation 7. The TMDCs have demonstrated good chemical stability and robust light-matter interactions, further improving their efficacy in light-driven processes. Despite these strengths, TMDCs are facing plausible limitations in photocatalytic hydrogen generation due to high exciton recombination and strong electrostatic coulomb interactions, leading to inefficient photogenerated charge carriers 8.
In comparison with conventional TMDCs, Janus TMDCs include different chalcogen atoms, breaking the out-of-plane symmetry 9. Thus, the Janus TMDCs will create out-of-plane dipoles between the top and bottom layers that could generate an intrinsic electric field and enhance the photocatalytic performance in Janus-TMDCs 10, 11. Based on the above reasons, the heterobilayers combining TMDCs and their Janus materials that have been successfully synthesized 12, 13 could benefit photocatalytic water splitting. In monolayer 2D material, both the hydrogen evolution reaction (HER) and the oxidation evolution reaction (OER) occur on the same surface of the 2D material, resulting in a higher recombination rate of photogenerated charge carriers 14, 15. A promising approach to overcome this limitation is the use of van der Waals (vdW) heterobilayers composed of two semiconductor layers.
These heterobilayers exhibit three primary types of heterostructure alignments: Type-I, Type-II, and Z-scheme, distinguished by their electronic band alignments and charge carrier dynamics. In the Type-I configuration, the conduction band minimum (CBM) and valence band maximum (VBM) of one semiconductor are entirely enclosed within the bandgap of the other 16. This alignment facilitates the transfer of both electrons and holes to a single material, making it suitable for applications requiring charge recombination, such as light emission.17 However, it is less effective for photocatalysis due to its limited ability to separate charge carriers. In contrast, the Type-II and Z-scheme heterobilayerss offer superior solutions. These configurations distribute the VBM and CBM across both semiconductor layers 18. In particular, the Z-scheme junction features a suitable band alignment and a wide directional built-in electric field. As a result, photogenerated charge carriers with a large redox potential are well-preserved and spatially separated within the Z-scheme heterobilayers. Combining two narrow band gap semiconductors into a Z-scheme configuration could solve the issue of the Type-II configuration, in which the redox ability and light absorption region may not be obtained simultaneously 19, 20, 18. The Z-scheme configuration allows HER and OER activity on separate surfaces 21.
@@
On the other hand, the heterobilayer with the Janus structure generates an intrinsic electric field, which not only enhances the photocatalytic performance but also accelerates the reaction kinetics by forming synergistic surface properties that allow selective adsorption of water molecules on each surface 22, 23. Moreover, the configuration of the heterobilayer can determine whether photogenerated carriers follow a Z-scheme or a traditional Type-II pathway 23, 24. However, the diversity of possible Janus-TMDC combinations in heterobilayers poses a significant challenge in determining the optimal configuration for photocatalytic water splitting. In this study, we provide a theoretical framework for rapidly identifying effective Janus-TMDC pairs through a top-down approach: candidates are initially filtered based on band alignment, and further refinement follows as needed. Another key challenge is the intrinsic polarization of Janus materials, which uniquely affects their electrical mobility. However, this factor is often neglected in previous Janus or TMDC studies for photocatalytic 25, 1, 26. To address this, we have adopted Fröhlich interaction model 27, 25 to account for the mobility adjustments in Janus heterobilayers accurately.
@@ 2 Theoretical Approach and Computational Details
In this research, we employed DFT for calculations using the Quantum ESPRESSO package 28. The generalized gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional was utilized as implemented in the standard solid-state pseudopotentials (SSSP) library 29, 30. Van der Waals interactions were incorporated using the Grimme DFT-D2 method 31, 32. To address the underestimation by PBE in predicting band edge positions and band gaps, we used the Heyd-Scuseria-Ernzerhof (HSE) hybrid-functional with a mixing parameter 25% 33, 34. The kinetic cutoff energy for wavefunctions ($E_{\text{cut}}$) was set to 800 eV, and the Monkhorst-Pack k-point grid was set to $12\times 12\times 1$ for PBE and $6\times 6\times 1$ for HSE based on the convergence test. All structures were fully relaxed using the BFGS scheme, with convergence criteria for force and energy differences set to $0.25\times 10^{-3}$ eV/Å and $1\times 10^{-5}$ eV, respectively. To prevent interlayer interactions, a relatively large vacuum layer of 30 Å was considered along the $z$-direction. To examine the Gibbs free energy and the mobility of the hydrogen atom, we adopted a $3\times 3\times 1$ supercell with 54 atoms. For the deformation potential theory (DPT) carrier mobility calculation, a rectangle $\sqrt{3}\times 1\times 1$ supercell was used to obtain the effective mass of carriers along $x$- and $y$-directions. In the calculation of mobility driven by Fröhlich scattering, the Born effective charge serves as a crucial parameter. It is determined by incorporating long-range electrostatic interactions through density functional perturbation theory 35. The Wannier90 code and the VESTA software were used for pre-processing and post-processing 36, 37.
Figure 1: Visual representations of the vdW Janus-TMDC structure: (a) top view and (b) side view.
@@ 3.1 Geometric and Electronic Structures
The geometric structures of these 20 materials are designed based on the previously synthesized AA stacking order of the MoSSe-MoS2 structure 38, systematically replacing the metal and chalcogenide atoms, as shown in Figure 1. In addition to the AA-stacked mode, the heterobilayers of TMDC/Janus heterostructures also exist in the AB-stacked mode. However, Nguyen et al. 9, 12 showed that the linear optical properties and electronic band structures of the AA- and AB-stacking orders are almost identical, which suggests that the photocatalytic characteristics of these orders should be the same. Thus, in the present study, we focus on the AA-stacking for 20 heterobilayers. The optimized lattice parameters of these heterobilayers are listed in Table 1. We note that the structures of MoS2-TeMoSe and WS2-TeWSe did not converge during the optimization calculations in this study. This non-convergence might come from the significant differences in electronegativity between the S and Te atoms at the interface.
Table 1: Calculated band gap energies $E_{g}$, work function differences $\Delta\Phi$, alignment type, and effective band gaps $E_{\text{eff}}$ for various Janus-TMDC heterobilayers.
Sample $E_{g}$ $\Delta\Phi$ Alignment $E_{\text{eff}}$
(eV) (eV) Type (eV)
MoS2-SeMoTe 0.32 0.82 Type-II -
MoS2-TeMoSe - - - -
MoSe2-SeMoTe 1.04 0.84 Type-I 1.04
MoSe2-TeMoSe 0.48 -0.60 Z-scheme 1.08
MoS2-SMoSe 1.31 0.82 Type-I 1.31
MoS2-SeMoS 0.68 -0.64 Z-scheme 1.32
MoSe2-SMoSe 0.49 0.71 Z-scheme 1.20
MoSe2-SeMoS 1.11 -0.73 Type-I 1.11
WS2-SWSe 1.44 0.79 Type-I 1.44
WS2-SeWS 0.82 -0.66 Z-scheme 1.48
WSe2-SWSe 0.60 0.69 Z-scheme 1.29
WSe2-SeWS 1.21 -0.70 Type-I 1.21
WS2-SeWTe 0.41 0.93 Type-II -
WS2-TeWSe - - - -
WSe2-SeWTe 1.10 0.80 Type-I 1.10
WSe2-TeWSe 0.97 -0.63 Z-scheme 1.60
MoS2-SWSe 1.56 0.83 Type-II 0.73
MoS2-SeWS 0.95 -0.58 Z-scheme 1.53
WS2-SMoSe 0.94 0.79 Z-scheme 1.72
WS2-SeMoS 1.52 -0.71 Z-scheme 2.23
@@ 3.2 Photocatalytic characteristic
After completion of structural optimizations, we evaluated the potential of 18 materials as photocatalysts (see Table S1 in Supporting Information). To determine the electronic properties critical for photocatalytic activity, we performed partial energy band structure for 18 material combinations, in which the partial band structures show the contributions of individual layers in the heterobilayers to the electronic states near the band edges (see Figure S1, Supporting Information). By projecting the electronic states onto the atomic orbitals of the Janus and TMDC layers, we identified which layers predominantly contribute to the VBM and CBM. This approach allowed us to classify the materials into Type I band alignment, where both the VBM and CBM are contributed by the same layer, or Type II alignment and Z-scheme alignment, where they are contributed by different layers, based on the spatial localization of their VBM and CBM.
@@
To be considered as a photocatalyst for water splitting, the theoretical band gap should be larger than 1.23 eV to overcome the reduction and oxidation potential of H2O 1, 40. However, due to the difference in electrostatic potential between the layers contributing to the VBM and CBM, even materials with smaller band gaps can act as photocatalysts. As shown in Table S1 (Supporting Information), the materials with Type-I band alignment, including MoSe2-SeMoTe, WSe2-SeWS, and WSe2-SeWTe, have band gaps less than 1.23 eV. Thus, they are unsuitable for water-splitting applications.
@@
For the remaining materials, we calculated the electrostatic potential curves, as shown in Figure S2 in Supporting Information. The calculations include a dipole correction to suppress artificial dipole fields by introducing an additional ramp-shaped potential 39. The intrinsic electric field generated by the Janus layer in the heterobilayers results in a significant potential difference between the two layers, leading to a difference in work function ($\Delta\Phi$) as calculated and presented in Table 1. Note that $\Delta\Phi$ is determined by taking the difference between the work functions of the TMDC and Janus layers, following this equation 8:
$\Delta\Phi=V_{\text{TMDC}}-V_{\text{Janus}},$ (1)
where $V_{\text{TMDC}}$ and $V_{\text{Janus}}$ are the vacuum energies of the TMDC and Janus layers in the heterobilayers, respectively. The values of $V_{\text{TMDC}}$ and $V_{\text{Janus}}$ of each heterobilayer are shown in Figure S2 on the Supporting Information.
@@
The band alignment plays a critical role in determining the photocatalytic properties of Janus-TMDC heterobilayers. When the work function difference between the layers, denoted as $\Delta\Phi$, is positive ($\Delta\Phi>0$), this indicates that the vacuum level of the TMDC layer is higher than that of the Janus layer. In such a configuration, the TMDC layer predominantly contributes to the VBM, while the Janus layer largely influences the CBM. This arrangement leads to a Z-scheme band alignment, which effectively broadens the separation between the CBM and VBM, enhancing charge carrier separation and improving photocatalytic activity (e.g., WSe2-SWSe and WS2-SMoSe, as shown in Figure 2) 19, 20, 18. Conversely, when $\Delta\Phi$ is negative ($\Delta\Phi<0$), meaning the vacuum level of the Janus layer is higher than that of the TMDC layer, the Janus layer mainly contributes to the VBM, and the TMDC layer to the CBM. This configuration also results in a Z-scheme alignment (e.g., WSe2-TeWSe, as shown in Figure 2).
@@
In cases where $\Delta\Phi$ is negligible or does not meet the conditions above, the system exhibits a Type-II band alignment. The effective band gap in a Z-scheme configuration can be expressed as $E_{\text{eff}}=E_{g}+\Delta\Phi$, where $E_{g}$ represents the intrinsic band gap, and $\Delta\Phi$ reflects the built-in potential difference at the heterobilayers interface. Materials with $E_{\text{eff}}>1.23$ eV are considered promising candidates for photocatalytic applications, and their performance is further evaluated through calculations of STH efficiency. The $E_{\text{eff}}$ values for the materials studied are summarized in Table 1. We note that the missing $E_{\text{eff}}$ values of WS2-SeWTe and MoS2-SeMoTe are due to the larger differences in vacuum potential, which result in CBM being positioned below the OER potential and VBM being above the HER potential. This placement makes the electron and hole at CBM and VBM recombination before doing HER and OER processes. As a result, these materials are unsuitable for water splitting.
3.3 Intrinsic strategies for band alignment control in Janus/TMDC heterobilayers
Designing heterobilayers of Janus and TMDC materials necessitates the control of interlayer electric fields to tailor band alignments for specific functional applications. Our findings indicate that manipulating these electric fields is key to achieving the desired type of band alignment. To investigate this, we calculate the electric field of the Janus/TMDC heterostructure through the electrostatic potential along the $z$-direction, as shown in Figure S2 (Supporting Information). The potential difference between the two chalcogenide layers inside $\Delta V_{\text{inner}}$ and outside $\Delta V_{\text{outer}}$ heterobilayers are defined as follows:
$\Delta V_{\text{inner}}=V_{\text{TMDC-inner}}-V_{\text{Janus-inner}},$ (2)
and
$\Delta V_{\text{outer}}=V_{\text{TMDC-outer}}-V_{\text{Janus-outer}},$ (3)
where $V_{\text{TMDC-inner}}$ and $V_{\text{Janus-inner}}$ are potential energies at the top chalcogenide of the TMDC layer and bottom chalcogenide of the Janus layer (see Figure 1 (b)), respectively. $V_{\text{TMDC-outer}}$ and $V_{\text{Janus-outer}}$ are potential energies at the bottom chalcogenide of the TMDC layer and top chalcogenide of the Janus layer, respectively.
Figure 3: Electrostatic potential differences across interface and vacuum regions in Janus and TMDC sites for Type-I, Type-II, and Z-Scheme alignments.
As illustrated in Figure 3 and Figure S3 (Supporting Information), materials with minimal internal electric fields at the interface tend to exhibit Type-I alignment, especially when the chalcogenide atoms at the interface are identical (i.e., $\Delta V_{\text{inner}}\sim 0$)-with the notable exception of MoS2–SWSe. In contrast, significant internal electric fields, driven by disparities in electronegativity between the contacting chalcogenide atoms, are strongly associated with Type-II and Z-scheme alignments. This means that a strong electric field at the Janus/TMDC interface can separate the CBM and VBM between the Janus and TMDC layers. Recent studies have highlighted that transitions between Type-I, Type-II, and Z-scheme alignments can be induced using external methods, such as applying strain, introducing external electric fields, or modifying interfacial composition via chemical functionalization 41, 10. For example, strain engineering has been shown to shift band alignments in ZrIBr/MgClBr and MoSTe/WSTe heterobilayers from Type-I to Type-II or Z-scheme, thereby enhancing their suitability for photovoltaic and catalytic applications 42.
Our findings emphasize the convenience of achieving these transitions intrinsically through electronegativity differences and atomic configurations at the interface. Notably, Type-II alignments generally align the internal electric field direction with the synthetic electric field of the heterostructure (i.e., $\Delta V_{\text{inner}}>0$ and $\Delta V_{\text{outer}}>0$), whereas Z-scheme alignments exhibit opposing directions (i.e., $\Delta V_{\text{inner}}>0$ and $\Delta V_{\text{outer}}<0$), enhancing charge separation and optimizing performance in photocatalysis (except MoSe2-TeMoSe and WSe2-TeWSe). When Janus and TMDC layers feature different metal atoms, Type-II or Z-scheme configurations are more likely, as the internal electric field aligns with the synthetic electric field of the outer chalcogenides. In contrast, heterobilayers with identical chalcogenide atoms often result in Type-I alignments, which typically confine photogenerated electrons and holes in the same layer, limiting charge separation and transport efficiency.
The mechanism behind the electric field tunes the band alignment from Type-I to Type-II or Z-scheme has been reported in previous studies, linking it to electron accumulation at the layer surface region 41, 10. Specifically, an increase in electron density occurs when the direction of the interfacial electric field shifts from pointing toward the contact region to pointing away from it. This shift leads to the stronger downward bending of the band levels in the surface region of heterobilayers with the electric field directed away from the contact area. Consequently, the band character of the heterobilayers transitions from Type-I to Type-II or Z-scheme as the direction of the interfacial electric field changes.
In photocatalytic applications, the objective is to leverage Z-scheme configurations, which mitigate the drawbacks of narrow intrinsic band gaps in monolayer TMDCs or Janus TMDCs while retaining strong light absorption and leading to the spatial separation of photogenerated electrons and holes. This spatial separation optimizes photocatalytic performance 18, 21. Type-I heterostructures typically replicate the narrow band gaps of their constituent layers, and Type-II configurations further narrow the overall band gap, reducing their effectiveness in photocatalysis. Given the intrinsic narrow band gaps of monolayer semiconductors like TMDCs, the water-splitting potential of Z-scheme heterostructures becomes crucial. This underscores the necessity of selecting suitable configurations during material screening to achieve optimal alignment for water-splitting photocatalysis. A strategic approach to predicting alignment types involves analyzing the electronegativity of the outermost chalcogenides to determine the direction of the overall electric field. When this field aligns with the internal electric field between layers, a Type-II alignment is likely; otherwise, the structure tends toward a Z-scheme configuration.
These insights offer a robust framework for designing Janus/TMDC heterobilayers with precisely controlled synthetic and internal electric fields, enabling the development of optimized materials. By strategically arranging atomic positions to manipulate electric fields, these heterobilayers can be tailored for specific applications in photocatalysis, light harvesting, and advanced optoelectronic devices 10, 41.